# SC-Bench: A Large-Scale Dataset for Smart Contract Auditing

### v1

@@ 3.2 Dataset Summary
Our dataset contains 15,975 ERC rule violations. Among them, 139 are introduced by the real-world programmers, while the remaining 15836 are injected by us. 39 errors made by real-world programmers cannot be replicated using the error-injection methods. These include 28 errors that violate the rule that transferring zero tokens must be treated the same as transferring non-zero tokens and 11 errors that violate the rule that the transfer() function must throw an exception if the sender does not have enough balance.
@@
Figures 4 and Figure 4 compare errors from two sources on their contract sizes and the average number of errors per contract On average, an originally violating contract contains 260.9 lines of source code, with a standard deviation of 198, whereas a contract with injected errors contains 477.5 lines of code, with a standard deviation of 385. Each originally violating contract contains 4.7 errors on average, and each modified contract contains 2.9 injected errors on average.
The security impact of the errors is tied to the methods used for error injection. A total of 3,645 errors arise from failures to emit events, resulting in contract activities going unlogged. These are considered to have a low-security impact. Among the 2230 high-security impact errors, a significant number are caused by missing required checks to verify sufficient privileges for performing specific actions. The lack of these checks can be easily exploited, leading to financial losses (e.g., stolen tokens, as illustrated in Figure 1). These errors are categorized as high-security impact. Additionally, failures to call required functions or incorrect updates to state variables can also have a high-security impact. The remaining 9961 errors cause contracts to behave unpredictably for users, although they may not directly result in financial loss, and are classified as having a medium-security impact. All these errors are injected through API, Value, Call, and Return methods.
Figure 4 shows the distribution of errors across different security impacts for the two sources. For errors identified through manual inspection, the proportions of high-impact, medium-impact, and low-impact errors are 14.8%, 43.0%, and 42.2%, respectively. For injected errors, the proportions are 51.3%, 24.5%, and 24.2%, respectively.
4 Evaluation
@@
Figure 5: The prompt template used to audit the contract. (For a concrete prompt, <ERC_type> could be ERC20, ERC721, and ERC1155, <code> is replaced with the whole contract code, and <ERC_content> is replaced with the description of one single rule or the whole ERC document.)
4.1 Methodology
For each auditing request, we ask GPT whether a contract violates any ERC rules, and if so, which function causes the violation. To instruct GPT on ERC rules, we adopt in-context learning by providing GPT ERC rules in addition to the contract under inspection, as shown in Figure 5. We use two methodologies to provide ERC rules. The first presents an ERC’s official document with all rules to GPT. After GPT returns its output, we compare both the GPT-generated violating rules and violating functions to the ground truth. We report when both are correct, when only the violating rule is correct, and when neither is correct.
The second method assumes oracle knowledge of which specific ERC rule(s) a contract violates and where the violation happens. So, it directly provides both to GPT and asks GPT a simple True-or-False question. A True result is correct (True Positive), while a False is wrong (False Negative). For this method, we manually select rules and violation code sites based on either the original contract violations or our injected errors.
4.2 Experimental Results
Full-Rule Prompting. As shown in Table 3, using full-rule prompting, GPT-4 successfully detects 139 errors (0.9% of the 15,975 errors), providing both the correct violated rules and the violating functions. For another 7 errors (0.04%), GPT-4 only reports the correct rule but fails to identify the correct location. For the remaining 15,830 errors (99%), GPT-4 does not provide any correct information.
We further separate the results between errors identified through manual inspection and those that are injected. For manually inspected errors, GPT-4 correctly reports both the violated rule and the violating function in 29% of the 139 errors. For injected errors, this proportion is 0.6%. The difference is probably due to that the injected errors are more challenging to identify than those introduced by programmers.
For errors in different security impacts, we notice GPT-4 has a very high detection rate when pinpointing manually injected errors with a high security impact, with a detection rate to be 80%. For errors in other security impacts, the detection rate ranges from 4% to 20.8%.
Table 3: Experimental Results for Full-Rule Prompting. ((x, y, z): x cases where both the reported rule and the reported violating function are correct, y cases where only the reported rule is correct, z cases where neither the rule nor the function is correct. “-”: all numbers are zeros.)
@@ 4.2 Experimental Results
Oracle Prompting. When using oracle prompting, GPT-4’s performance significantly improves, with the detection rate increasing from 0.9% to 22.9%. The improvement varies depending on the source of the errors. For manually inspected errors, the detection rate rises from 29% to 42.7%, while for injected errors, it increases from 0.6% to 22.8%. This boost in detection is mainly due to breaking down a complex task into smaller, more manageable tasks, allowing GPT-4 to focus on each one individually.
Implications. Our results show that simple prompting techniques even with oracle still have a large room for accuracy improvements. A major reason we suspect is the complexity of smart contracts and the original rule description, which prevents GPT-4 from properly detecting violations. Potential ways to improve accuracy include further narrowing down contract sizes (e.g., by slicing code) and improving the understandability of ERC rules (e.g., through specialized prompts).
Table 4: Experimental Results for Oracle Prompting. ((x, y): x true positives, and y false negatives. “-”: all numbers are zeros.)
@@
ERC1155 - - - - (4,27) (0,30) (0,0) (4,57) (4,57)
Total (3,18) (29,30) (30,35) (62,83) (745,1485) (2038,7923) (823, 2822) (3606, 12230) (3668,12313)
5 Discussion and Conclusion
@@
We present SC-Bench, the first dataset for smart contract auditing. SC-Bench contains 5377 real-world smart contracts and 15975 ERC violations, with two sources of violations: real-world errors and injected errors. Our evaluation of SC-Bench with GPT-4 shows that while ML-based techniques are promising, there is still huge room for improvement.
@@
Notably, SC-Bench has certain limitations. For example, we only include the three most important ERC standards and only inject one to three errors per contract. Adding more ERC standards and supporting more types of error injection is feasible — something we leave for future work. Another limitation is the imbalance of real-world violations and injected errors. As real-world violations are rarely reported, we believe the imbalance cannot be easily improved. One possible approach is to instruct ML models to create real-world-like errors.

### v2

@@ III-B Dataset Summary
Our dataset contains 15,975 ERC rule violations. Among them, 139 are introduced by the real-world programmers, while the remaining 15,836 are injected by us. 39 errors made by real-world programmers cannot be replicated using the error-injection methods. These include 28 errors that violate the rule that transferring zero tokens must be treated the same as transferring non-zero tokens and 11 errors that violate the rule that the transfer() function must throw an exception if the sender does not have enough balance.
@@
Figures 6 and Figure 6 compare errors from two sources on their contract sizes and the average number of errors per contract. On average, an originally violating contract contains 260.9 lines of source code, with a standard deviation of 198, whereas a contract with injected errors contains 477.5 lines of code, with a standard deviation of 385. Each originally violating contract contains 4.83 errors on average, and each modified contract contains 2.96 injected errors on average.
The security impact of the errors is tied to the methods used for error injection. A total of 3,645 errors arise from failures to emit events, resulting in contract activities going unlogged. These are considered to have a low-security impact. Among the 2,230 high-security impact errors, a significant number are caused by missing required checks to verify sufficient privileges for performing specific actions. The lack of these checks can be easily exploited, leading to financial losses (e.g., stolen tokens, as illustrated in Figure 1). These errors are categorized as having a high-security impact. Additionally, failures to call required functions or incorrect updates to state variables can also have a high-security impact. The remaining 9,961 errors cause contracts to behave unpredictably for users, although they may not directly result in financial loss, and are classified as having a medium-security impact. All these errors are injected through API, Value, Call, and Return methods.
Figure 6 shows the distribution of errors across different security impacts for the two sources. For errors identified through manual inspection, the proportions of high-impact, medium-impact, and low-impact errors are 9.4%, 29.2%, and 61.3%, respectively. For injected errors, the proportions are 14.1%, 62.9%, and 23.0%, respectively.
IV Evaluation
@@
Figure 7: The full-rule prompting template. (For a concrete prompt, <ERC_type> could be ERC20, ERC721, and ERC1155, <code> is replaced with the whole contract code, and <ERC_content> is replaced with the whole ERC document.)
1Whether the following smart contract <contract_name> violate ERC rule <rule> for function <function_sig>? Answer YES or NO.
2Code:"""
3<code>
4"""
Figure 8: The oracle prompting template. (For a concrete prompt, <contract_name>, <rule>, <function_sig>, and <code> are replaced with the name of the smart contract, the rule’s natural language description, the declaration of the function, and the code of the function and all its callees.)
1[
2 {
3 "rule": "approve: Allows _spender to withdraw from your account multiple times, up to the _value amount. If this function is called again it overwrites the current allowance with _value.",
4 "function": "approve function is missing in the contract"
5 },
6 {
7 "rule": "Approval: MUST trigger on any successful call to approve(address _spender, uint256 _value).",
8 "function": "Approval event is missing in the contract"
9 }
10]
Figure 9: The GPT response for the contract in Figure 1 with full-rule prompting.
TABLE III: Experimental results for full-rule prompting. ((x, y, z): x cases where both the reported rule and the reported violating function are correct, y cases where only the reported rule is correct, z cases where neither the rule nor the function is correct. “-”: all numbers are zeros.)
@@
TABLE IV: Experimental results for oracle prompting. ((x, y): x true positives, and y false negatives. “-”: all numbers are zeros.)
@@
ERC1155 - - - - (4,27) (0,30) - (4,57) (4,57)
Total (3,24) (29,19) (30,34) (62,77) (745,1485) (2038,7923) (823, 2822) (3606, 12230) (3668,12307)
IV-A Methodology
For each auditing request, we ask GPT whether a contract violates any ERC rules, and if so, which function causes the violation. To instruct GPT on ERC rules, we adopt in-context learning by providing GPT ERC rules in addition to the contract under inspection. We use two methodologies to provide ERC rules. The first presents an ERC’s official document with all rules to GPT, as shown in Figure 7. After GPT returns its output, we compare both the GPT-generated violating rules and violating functions to the ground truth. We report when both are correct, when only the violating rule is correct, and when neither is correct.
The second method assumes oracle knowledge of which specific ERC rule(s) a contract violates and the function where the violation happens. So, it directly provides both to GPT and asks GPT a simple True-or-False question, as shown by Figure 8. A True result is correct (True Positive), while a False is wrong (False Negative). For this method, we manually select rules and violation code sites based on either the original contract violations or our injected errors.
IV-B Experimental Results
Full-Rule Prompting. As shown in Table III, using full-rule prompting, GPT-4 successfully detects 144 errors (0.9% of the 15,975 errors), providing both the correct violated rules and the violating functions. For another 7 errors (0.04%), GPT-4 only reports the correct rule but fails to identify the correct location. For the remaining 15,824 errors (99%), GPT-4 does not provide any correct information.
For instance, Figure 9 illustrates GPT-4’s response when analyzing the contract in Figure 1 using full-rule prompting. GPT-4 accurately identifies two rule violations and the functions responsible for them. Since the contract does not include the function approve(address _spender, uint256 _value), it is considered violating both that the function should overwrite the current allowance with _value and that the function should emit the Approval event. However, GPT-4 fails to recognize that the contract fails to ensure the transferFrom() function properly validates its caller’s privileges, missing a critical violation with a high security impact. In another example, when asked to audit the contract in Figure 3, GPT-4 fails to detect that the balanceOf(address account) function does not return the expected value as required by ERC20.
We further separate the results between errors identified through manual inspection and those that are injected. For manually inspected errors, GPT-4 correctly reports both the violated rule and the violating function in 30% of the 139 errors. For injected errors, this proportion is 0.6%. The difference is probably due to that the injected errors are more challenging to identify than those introduced by programmers.
For errors in different security impacts, we notice GPT-4 has a very high detection rate when pinpointing manually introduced errors with a high security impact, with a detection rate to be 59.2%. For errors in other security impacts, the detection rate ranges from 0.13% to 43.8%.
Oracle Prompting. Table IV shows the results with oracle prompting. Under this setting, GPT-4’s performance significantly improves, with the detection rate increasing from 0.9% to 22.9%. The improvement varies depending on the source of the errors. For manually inspected errors, the detection rate rises from 30% to 44.6%, while for injected errors, it increases from 0.6% to 22.8%. This boost in detection is mainly due to breaking down a complex task into smaller, more manageable tasks, allowing GPT-4 to focus on each one individually. Those results show that simple prompting techniques even with oracle still have a large room for accuracy improvements.
For example, when presenting both the source code of the transferFrom(address from, address to, uint256 amount) function and its callee from lines 6 to 17 in Figure 1, along with the rule description stating that the function “should throw unless the _from account has deliberately authorized the sender of the message via some mechanism,” and asking GPT-4 whether the rule is violated, GPT-4 correctly answers yes. In another example, when showing the balanceOf() function from lines 4 to 7 in Figure 3, along with the rule that the function “returns the account balance of another account with address _owner,” and asking GPT-4 to analyze if the rule is violated, it correctly responds yes.
V Discussion and Conclusion
@@
We present SC-Bench, the first dataset for smart contract auditing. SC-Bench contains 5,377 real-world smart contracts and 15,975 ERC violations, with two sources of violations: real-world errors and injected errors. Our evaluation of SC-Bench with GPT-4 shows that while ML-based techniques are promising, there is still huge room for improvement.
@@
Notably, SC-Bench has certain limitations. For example, we only include the three most important ERC standards and only inject one to three errors per contract. Both adding more ERC standards and supporting more types of error injection are feasible — something we leave for future work. Another limitation is the imbalance of real-world violations and injected errors. As real-world violations are rarely reported, we believe the imbalance cannot be easily improved. One possible approach is to instruct ML models to create real-world-like errors.